OpenStudy (anonymous):

The domain bounded by the surface of a paraboloid z=2-x^2-y^2 and that of a cone z^2=x^2+y^2 is given by D={x,y,z:x^2+y^2<=1, sqrt(x^2+y^2)<=z<=2-x^2-y^2} can someone explain how to get the upper limits for x,y,z please

OpenStudy (turingtest):

change this to cylindrical coordinates

OpenStudy (anonymous):

I need to do this triple intergral intergration and just need to understand how to get the upper limits for x,y,z in order to intergrate

OpenStudy (anonymous):

can you explain this please if you are able to?

OpenStudy (turingtest):

in rectangular coordinates your bounds are sqrt(x^2+y^2)<=z<=2-x^2-y^2 0<=y<=sqrt(1-x^2) 0<=x<=1 but that is an ugly integral, I strongly recommend switching to cylindrical coordinates

OpenStudy (anonymous):

ok, thanks for the the bounds but I would just appreciate it if you could explain how you came up with the bounds, sorry to be so thick.

OpenStudy (turingtest):

sorry I meant sqrt(x^2+y^2)<=z<=2-x^2-y^2 -sqrt(1-x^2)<=y<=sqrt(1-x^2) 0<=x<=1 you get the bounds for z directly from the area: sqrt(x^2+y^2)<=z<=2-x^2-y^2 you are also given that it is over the region x^2+y^2=1 solve this for one variable: -sqrt(1-x^2)<=y<=sqrt(1-x^2) since the argument under the square root cannot be negative we see that x cannot be greater than 1,hence the bounds on x are 0<=x<=1

OpenStudy (anonymous):

thank you sooo much for that explanation

OpenStudy (anonymous):

now how would you change it to polar coordinates?

OpenStudy (turingtest):

cylindrical coordinates: \[x=r\cos\theta\]\[y=r\sin\theta\]\[z=z\]\[dA=rdzdrd\theta\]use the symmetry of the object to find the bounds on r and theta

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you!

OpenStudy (turingtest):

welcome!